Question: An investor has a total of $25,000 deposited in three different accounts, which earn annual interest rates of 8%, 6%, and 4%. The amount deposited in the 8% account is twice the amount in the 6% account. If the three accounts earn total interest of $1,520, how much money is deposited in each account.
Note: I have not attempted this problem because it confuses me and we are doing two of the four boxed problems but I don't see how to work it here.

Answer by mananth(16946)

An investor has a total of $25,000 deposited in three different accounts, which earn annual interest rates of 8%, 6%, and 4%. The amount deposited in the 8% account is twice the amount in the 6% account. If the three accounts earn total interest of $1,520, how much money is deposited in each account.
let amount invested at 6% be x
8%-----------2x
Balance at 4%
4%-----------(25000-3x)
6%*x +8%*2x +4%(25000-3x)=1520
multiply equation by 100
6x+16x+100,000-12x=152000
10x=152000-100000
10x=52000
x= 5,200-----------6%
2x=10,400----------8%
8,400--------------4%
Check
6%*5,200 +8%*10,400 +4%*9,400= 1520